OpenStudy (anonymous):

The square of a number is equal to 10 less than 7 times that number. What are the two possible solutions? Which of the following equations is used in the process of solving this problem?

OpenStudy (anonymous):

Let number be X. Square of number = x^2 equal to 10 less than 7 times the number x^2 = 7x - 10 x^2 - 7x + 10 = 0 solve for x?

OpenStudy (anonymous):

thanks.